Update for the Week Ending May 12, 2017

Site Remediation Completed
Site Restoration Underway

On-going activities:

  • Continued the application of water for dust control;
  • Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip and straw filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site);
  • Continued final grading of select fill in preparation for top soil placement;
  • Initiated the import and stockpiling of top soil on Site;
  • Initiated the shipment of vacuum boxes of sediment from the WTP storage tank decontamination to Bayshore Soil Management (BSM) for treatment and disposal;
  • Initiated the installation of Site fencing;
  • Continued decontaminating and demobilizing equipment from the Site; and
  • Continued Site perimeter preparation for new fencing and damaged sidewalk replacement.
Recently Completed:

  • Completed shallow excavation in the western areas of the Site;
  • Completed lining and loading trucks and shipping impacted soil to Bayshore Soil Management (BSM) for treatment and disposal:
    • 5 trucks; 2 vacuum boxes of sediment from the WTP storage tanks;
    • 1,194.74 tons shipped last week; and
    • 65,136.58 tons shipped to BSM to date.
  • Completed pumping from the decontamination pad and water treatment plant (WTP) component decontamination to the last remaining WTP storage tank;
  • Completed the decontamination of the WTP components;
  • Completed the import of select fill backfill to the Site;
  • Completed backfilling in the deep and shallow excavation areas with select fill;
  • Completed compaction of the backfill lifts; and
  • Completed compaction testing of the backfill lifts.

Upcoming activities:

  • Import and place topsoil;
  • Hydroseed the Site;
  • Replace damaged sidewalks along N. 12th Street, Kent Avenue, and N. 11th Street;
  • Install fencing around the 50 Kent Avenue property; and
  • Complete the demobilization of equipment from the Site and support areas along N. 12th Street.

Community Air Monitoring Plan Summary

Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.

There were no 15-minute Action Limit exceedances for Total Volatile Organic Compounds (TVOC), Particulate Matter (PM-10) [dust], or odors.

If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.

Photos of Activity


Offloading select fill backfill to the Site.


Loading lined trucks with impacted soil for shipment to BSM for treatment and disposal.


Loading vacuum boxes of impacted sediment for shipment to BSM for treatment and disposal.


Grading select fill backfill in the eastern shallow excavation area.


Compaction testing select fill backfill lifts in the western shallow excavation area.


Surveying backfill grades on Site.


Offloading and stockpiling top soil on Site.